OpenStudy (anonymous):

What is the solution to the system of equations represented by these two lines? A. (3, 1) B. (3, 0) C. (1, 3) D. (0, 6)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

where are the lines dear? we need those!

OpenStudy (anonymous):

Ok :) one sec.

OpenStudy (mitu12):

yesss

OpenStudy (anonymous):

what grade

OpenStudy (mitu12):

which point does the line intersect

OpenStudy (anonymous):

I've got it misty1212

OpenStudy (anonymous):

@misty1212

OpenStudy (mitu12):

the ordered pair where the line intersects is your answer

OpenStudy (anonymous):

So it's c?

OpenStudy (anonymous):

@mitu12

OpenStudy (mitu12):

it would be A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you help with a few more?

OpenStudy (mitu12):

sure

OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=-2x+7 y=1/3x

OpenStudy (mitu12):

\[-2x+7=1/3x\] which two like terms can you combine

OpenStudy (anonymous):

2x and 3x

OpenStudy (mitu12):

subtract either number from both sides

OpenStudy (anonymous):

-1 and 2

OpenStudy (mitu12):

if you add 2x to 1/3x what do you get

OpenStudy (anonymous):

5x and 1

OpenStudy (mitu12):

i dont think im sure on how to do this problem sorry

OpenStudy (anonymous):

ok............